# Label Distribution Learning using the Squared Neural Family on the Probability Simplex

### v1

@@ Appendix B Label Distribution Prediction Results
Table 6: The comparison of label distribution prediction performance measured by Clark distance $\downarrow$.
Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
Movie 1.0116 $\spadesuit$ 1.0205 $\spadesuit$ 1.0588 $\bullet$ 1.0055 $\spadesuit$ 1.4036 $\bullet$ 0.9796 $\circ$ 1.1795 $\bullet$ 0.9970 $\circ$ 1.0154 +0.0357
Natural_Scene 6.7001 $\spadesuit$ 4.5465 $\circ$ 6.9670 $\bullet$ 6.6520 $\circ$ 6.6519 $\spadesuit$ 6.5907 $\circ$ 6.5539 $\circ$ 6.6785 $\circ$ 6.7202 +2.1737
SBU_3DFE 0.8713 $\bullet$ 0.8306 $\bullet$ 0.9055 $\bullet$ 0.7827 $\bullet$ 0.8035 $\bullet$ 0.7638 $\bullet$ 0.8580 $\bullet$ 0.8485 $\bullet$ 0.7347 -0.0291
SJAFFE 0.8604 $\bullet$ 0.7148 $\spadesuit$ 0.9001 $\bullet$ 0.7326 $\spadesuit$ 0.6820 $\circ$ 0.6188 $\circ$ 0.6648 $\circ$ 0.6421 $\circ$ 0.7473 +0.1285
Yeast_alpha 0.7237 $\bullet$ 0.7388 $\bullet$ 0.7327 $\bullet$ 0.6818 $\circ$ 0.7051 $\spadesuit$ 0.6802 $\circ$ 0.6911 $\circ$ 0.6819 $\circ$ 0.7128 +0.0325
Yeast_cdc 0.6794 $\bullet$ 0.7001 $\bullet$ 0.6733 $\spadesuit$ 0.6473 $\circ$ 0.6524 $\circ$ 0.6472 $\circ$ 0.6540 $\circ$ 0.6475 $\circ$ 0.6673 +0.0202
Yeast_cold 0.2487 $\bullet$ 0.2550 $\bullet$ 0.2515 $\bullet$ 0.2403 $\spadesuit$ 0.2513 $\bullet$ 0.2424 $\spadesuit$ 0.2558 $\bullet$ 0.2403 $\spadesuit$ 0.2430 +0.0027
Yeast_diau 0.4501 $\bullet$ 0.4476 $\spadesuit$ 0.4671 $\bullet$ 0.4308 $\circ$ 0.4367 $\spadesuit$ 0.4300 $\circ$ 0.4804 $\bullet$ 0.4308 $\circ$ 0.4449 +0.0149
Yeast_dtt 0.1796 $\bullet$ 0.1789 $\bullet$ 0.1732 $\spadesuit$ 0.1689 $\circ$ 0.1750 $\spadesuit$ 0.1711 $\spadesuit$ 0.1832 $\spadesuit$ 0.1689 $\circ$ 0.1731 +0.0043
Yeast_elu 0.6236 $\bullet$ 0.6281 $\bullet$ 0.6084 $\spadesuit$ 0.5831 $\circ$ 0.5876 $\circ$ 0.5820 $\circ$ 0.5905 $\circ$ 0.5832 $\circ$ 0.6116 +0.0296
Yeast_heat 0.3780 $\bullet$ 0.3834 $\bullet$ 0.3739 $\bullet$ 0.3641 $\circ$ 0.3688 $\spadesuit$ 0.3626 $\circ$ 0.3763 $\bullet$ 0.3642 $\circ$ 0.3670 +0.0044
Yeast_spo 0.5199 $\bullet$ 0.5489 $\bullet$ 0.5223 $\bullet$ 0.5131 $\spadesuit$ 0.5346 $\bullet$ 0.5245 $\bullet$ 0.5218 $\bullet$ 0.5129 $\spadesuit$ 0.5143 +0.0014
Yeast_spo5 0.2856 $\bullet$ 0.2945 $\bullet$ 0.2847 $\bullet$ 0.2826 $\spadesuit$ 0.3082 $\bullet$ 0.3003 $\bullet$ 0.3344 $\bullet$ 0.2824 $\spadesuit$ 0.2820 -0.0004
Yeast_spoem 0.1828 $\bullet$ 0.1878 $\bullet$ 0.1807 $\spadesuit$ 0.1797 $\spadesuit$ 0.1934 $\bullet$ 0.2149 $\bullet$ 0.4967 $\bullet$ 0.1791 $\spadesuit$ 0.1792 +0.0001
Table 7: The comparison of label distribution prediction performance measured by Canberra metric $\downarrow$.
Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
Movie 0.1011 $\bullet$ 0.1093 $\bullet$ 0.1113 $\bullet$ 0.1002 $\bullet$ 0.2357 $\bullet$ 0.0975 $\spadesuit$ 0.1380 $\bullet$ 0.0981 $\spadesuit$ 0.0984 +0.0009
Natural_Scene 0.7608 $\spadesuit$ 1.0936 $\bullet$ 1.0466 $\bullet$ 0.8277 $\spadesuit$ 0.7832 $\spadesuit$ 0.7608 $\spadesuit$ 1.0478 $\bullet$ 0.7445 $\spadesuit$ 0.7704 +0.0259
SBU_3DFE 0.0774 $\bullet$ 0.0803 $\bullet$ 0.0860 $\bullet$ 0.0629 $\bullet$ 0.0651 $\bullet$ 0.0620 $\bullet$ 0.0776 $\bullet$ 0.0730 $\bullet$ 0.0578 -0.0042
SJAFFE 0.0701 $\bullet$ 0.0534 $\spadesuit$ 0.0740 $\bullet$ 0.0505 $\spadesuit$ 0.0446 $\circ$ 0.0375 $\circ$ 0.0419 $\circ$ 0.0391 $\circ$ 0.0523 +0.0148
Yeast_alpha 0.0060 $\bullet$ 0.0063 $\bullet$ 0.0061 $\bullet$ 0.0055 $\circ$ 0.0058 $\spadesuit$ 0.0055 $\circ$ 0.0056 $\circ$ 0.0055 $\circ$ 0.0059 +0.0004
Yeast_cdc 0.0075 $\bullet$ 0.0079 $\bullet$ 0.0076 $\bullet$ 0.0070 $\circ$ 0.0071 $\circ$ 0.0070 $\circ$ 0.0071 $\circ$ 0.0070 $\circ$ 0.0073 +0.0004
Yeast_cold 0.0128 $\bullet$ 0.0136 $\bullet$ 0.0132 $\bullet$ 0.0122 $\circ$ 0.0132 $\bullet$ 0.0123 $\spadesuit$ 0.0137 $\bullet$ 0.0122 $\circ$ 0.0124 +0.0002
Yeast_diau 0.0142 $\bullet$ 0.0145 $\spadesuit$ 0.0151 $\bullet$ 0.0131 $\circ$ 0.0138 $\spadesuit$ 0.0132 $\circ$ 0.0165 $\bullet$ 0.0131 $\circ$ 0.0138 +0.0007
Yeast_dtt 0.0069 $\bullet$ 0.0070 $\bullet$ 0.0065 $\spadesuit$ 0.0063 $\circ$ 0.0068 $\spadesuit$ 0.0064 $\spadesuit$ 0.0072 $\spadesuit$ 0.0063 $\circ$ 0.0065 +0.0002
Yeast_elu 0.0069 $\bullet$ 0.0071 $\bullet$ 0.0066 $\spadesuit$ 0.0062 $\circ$ 0.0063 $\circ$ 0.0062 $\circ$ 0.0063 $\circ$ 0.0062 $\circ$ 0.0066 +0.0005
Yeast_heat 0.0135 $\bullet$ 0.0139 $\bullet$ 0.0133 $\bullet$ 0.0126 $\circ$ 0.0130 $\spadesuit$ 0.0125 $\circ$ 0.0133 $\spadesuit$ 0.0126 $\circ$ 0.0129 +0.0004
Yeast_spo 0.0251 $\spadesuit$ 0.0289 $\bullet$ 0.0258 $\bullet$ 0.0246 $\circ$ 0.0271 $\bullet$ 0.0259 $\bullet$ 0.0257 $\bullet$ 0.0246 $\circ$ 0.0251 +0.0005
Yeast_spo5 0.0298 $\bullet$ 0.0326 $\bullet$ 0.0297 $\bullet$ 0.0292 $\spadesuit$ 0.0351 $\bullet$ 0.0338 $\bullet$ 0.0409 $\bullet$ 0.0292 $\spadesuit$ 0.0290 -0.0002
Yeast_spoem 0.0253 $\bullet$ 0.0274 $\bullet$ 0.0247 $\spadesuit$ 0.0245 $\spadesuit$ 0.0301 $\bullet$ 0.0346 $\bullet$ 0.1475 $\bullet$ 0.0245 $\spadesuit$ 0.0243 -0.0002
Table 8: The comparison of label distribution prediction performance measured by Kullback-Leibler divergence $\downarrow$.
Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
Movie 0.8324 $\spadesuit$ 0.8294 $\bullet$ 0.8196 $\bullet$ 0.8336 $\spadesuit$ 0.7572 $\bullet$ 0.8380 $\circ$ 0.7953 $\bullet$ 0.8353 $\circ$ 0.8333 -0.0048
Natural_Scene 0.5549 $\spadesuit$ 0.5629 $\circ$ 0.4153 $\bullet$ 0.5491 $\spadesuit$ 0.5536 $\spadesuit$ 0.5566 $\spadesuit$ 0.5294 $\spadesuit$ 0.5595 $\circ$ 0.5343 -0.0286
SBU_3DFE 0.8436 $\bullet$ 0.8484 $\bullet$ 0.8380 $\bullet$ 0.8600 $\bullet$ 0.8559 $\bullet$ 0.8624 $\bullet$ 0.8456 $\bullet$ 0.8477 $\bullet$ 0.8685 +0.0061
SJAFFE 0.8523 $\bullet$ 0.8768 $\spadesuit$ 0.8466 $\bullet$ 0.8786 $\spadesuit$ 0.8865 $\circ$ 0.8959 $\circ$ 0.8876 $\circ$ 0.8928 $\circ$ 0.8729 -0.0229
Yeast_alpha 0.9600 $\bullet$ 0.9592 $\bullet$ 0.9595 $\bullet$ 0.9624 $\circ$ 0.9611 $\spadesuit$ 0.9624 $\circ$ 0.9618 $\circ$ 0.9624 $\circ$ 0.9606 -0.0018
Yeast_cdc 0.9552 $\bullet$ 0.9539 $\bullet$ 0.9557 $\spadesuit$ 0.9574 $\circ$ 0.9571 $\circ$ 0.9574 $\circ$ 0.9570 $\circ$ 0.9574 $\circ$ 0.9561 -0.0014
Yeast_cold 0.9387 $\bullet$ 0.9371 $\bullet$ 0.9379 $\bullet$ 0.9408 $\spadesuit$ 0.9380 $\bullet$ 0.9403 $\spadesuit$ 0.9368 $\bullet$ 0.9408 $\spadesuit$ 0.9401 -0.0007
Yeast_diau 0.9375 $\bullet$ 0.9379 $\spadesuit$ 0.9350 $\bullet$ 0.9403 $\circ$ 0.9394 $\spadesuit$ 0.9404 $\circ$ 0.9334 $\bullet$ 0.9403 $\circ$ 0.9383 -0.0021
Yeast_dtt 0.9556 $\bullet$ 0.9559 $\bullet$ 0.9572 $\spadesuit$ 0.9583 $\circ$ 0.9568 $\spadesuit$ 0.9578 $\spadesuit$ 0.9547 $\spadesuit$ 0.9583 $\circ$ 0.9573 -0.0011
Yeast_elu 0.9560 $\bullet$ 0.9557 $\bullet$ 0.9571 $\spadesuit$ 0.9589 $\circ$ 0.9585 $\circ$ 0.9589 $\circ$ 0.9583 $\circ$ 0.9589 $\circ$ 0.9568 -0.0021
Yeast_heat 0.9379 $\bullet$ 0.9371 $\bullet$ 0.9386 $\bullet$ 0.9402 $\circ$ 0.9395 $\spadesuit$ 0.9405 $\circ$ 0.9382 $\bullet$ 0.9402 $\circ$ 0.9397 -0.0008
Yeast_spo 0.9144 $\bullet$ 0.9094 $\bullet$ 0.9140 $\bullet$ 0.9155 $\spadesuit$ 0.9118 $\bullet$ 0.9135 $\bullet$ 0.9140 $\bullet$ 0.9156 $\spadesuit$ 0.9153 -0.0002
Yeast_spo5 0.9077 $\bullet$ 0.9048 $\bullet$ 0.9080 $\bullet$ 0.9087 $\spadesuit$ 0.9002 $\bullet$ 0.9027 $\bullet$ 0.8917 $\bullet$ 0.9088 $\spadesuit$ 0.9090 +0.0002
Yeast_spoem 0.9116 $\bullet$ 0.9093 $\bullet$ 0.9126 $\spadesuit$ 0.9132 $\spadesuit$ 0.9068 $\bullet$ 0.8957 $\bullet$ 0.7685 $\bullet$ 0.9135 $\spadesuit$ 0.9135 +0.0000
Table 9: The comparison of label distribution prediction performance measured by Intersection $\uparrow$.
@@ Appendix C Parameter Sensitivity Study
On the Movie and SBU_3DFE datasets, we take turns to study the sensitivity of SNEFY-LDL with regard to the four hyperparameters: $n$ and $m$, as well as the batch size and epoch number used for training, by varying the studied hyperparameter in a predefined range and fixing the remaining three as default values at each turn. Figure 1 plots the performance change of SNEFY-LDL measured by Cosine coefficient with the varying values of the four hyperparameters. From Figure 1, we can find that the performance of SNEFY-LDL remains relatively stable when the hyperparameters vary in proper ranges.
@@
Figure 1: Parameter sensitivity.

### v2

@@ B.3 Ensemble Learning
10:  return the label distribution prediction performance scores measured by Cheby, Clark, Canb, KL, Cos and Inter.
Algorithm A4 Ensemble Learning with SNEFY-LDL
[…]
Parameters: The number of samples for training base learners $N_{\text{sample}}=50$ and the number of base learners $N_{\text{base}}=25$.
Output: The label distribution prediction performance scores measured by Cheby, Clark, Canb, KL, Cos and Inter on the test set $\mathcal{S}^{\text{te}}$.
1:  for each iteration $i\in\{1,\cdots,N_{\text{base}}\}$ do
2:     Randomly select $N_{\text{sample}}$ samples from the training set $\mathcal{S}^{\text{tr}}$;
3:     Train a LDL base learner $B_{i}$ with the selected samples;
4:  end for
5:  Train a SNEFY-LDL model with the training set $\mathcal{S}^{\text{tr}}$;
6:  for each sample $\bm{x}\in\mathcal{S}^{\text{te}}$ do
7:     for each iteration $i\in\{1,\cdots,N_{\text{base}}\}$ do
8:        Predict $\bm{x}$’s label distribution vector with base learner $B_{i}$ as ~$\tilde{\bm{\ell}}^{(i)}_{\bm{x}}$;
9:        Evaluate the probability density value $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})/d\bm{\ell}$ at =~$\bm{\ell}=\tilde{\bm{\ell}}^{(i)}_{\bm{x}}$ as $p(\tilde{\bm{\ell}}^{(i)}_{\bm{x}}|\bm{x};\bm{V},\bm{\Theta})$ with Eq. (6) and the trained SNEFY-LDL model;
10:     end for
11:     Predict $\bm{x}$’s label distribution vector ~$\tilde{\bm{\ell}}_{\bm{x}}$ as the weighted average of ~$\tilde{\bm{\ell}}^{(i)}_{\bm{x}}$, i.e., ~$\tilde{\bm{\ell}}_{\bm{x}}=\frac{\sum_{i=1}^{N_{\text{base}}}p(\tilde{\bm{\ell}}^{(i)}_{\bm{x}}|\bm{x};\bm{V},\bm{\Theta})\tilde{\bm{\ell}}^{(i)}_{\bm{x}}}{\sum_{i=1}^{N_{\text{base}}}p(\tilde{\bm{\ell}}^{(i)}_{\bm{x}}|\bm{x};\bm{V},\bm{\Theta})}$;
12:  end for
13:  Evaluate the performance of the ensembled label distribution predictions ~$\tilde{\bm{\ell}}^{(i)}_{\bm{x}}$ on the test set $\mathcal{S}^{\text{te}}$ (as an average over all test samples) with the metrics of Cheby, Clark, Canb, KL, Cos and Inter defined in Eq. (22);
14:  return the label distribution prediction performance scores measured by Cheby, Clark, Canb, KL, Cos and Inter.
@@ Appendix C Parameter Sensitivity Study
By choosing the Natural Scene dataset and the conformal prediction task, we take turns to study the sensitivity of SNEFY-LDL with regard to the four hyperparameters: $n$ and $m$, as well as the batch size and epoch number used for training, by varying the studied hyperparameter in a predefined range and fixing the remaining three as default values at each turn. For the conformal prediction task, the default values of $n$, $m$, batch size and epoch number are 64, 32, 64 and 100 respectively. Figure A1 plots the conformal prediction performance change of SNEFY-LDL measured by FSC with bin size equal to 2 and 4 when the values of hyperparameters vary in a range. From Figure A1, we can find that the performance of SNEFY-LDL remains relatively stable with the change of the four hyperparameters in most cases, except for the cases with $m=48$ and $96$ in Figure A1(b), as well as batch size $=32$ and $96$ in Figure A1(c), where the conformal prediction performance of SNEFY-LDL goes through a obvious drop.
@@
Figure A1: The sensitivity of SNEFY-LDL with regard to the four hyperparameters: $n$, $m$, batch size and epoch number.